If LN=6x-5, LM=x+7, and MN=3x+20, find MN.
1
Expert's answer
2020-10-04T11:53:24-0400

"LM+MN=LN"

"x+7+3x+20=6x-5"

"4x+27=6x-5"

"4x-6x=-27-5"

"-2x=-32"

"x=-32:(-2)"

"x=16"

Therefore

"MN=3x+20=3\u22c516+20=48+20=68"

Answer: "MN=68".